a At 37 °C, if the plasma pH is 7.40 and the plasma concentration of HCO3 is 12.0 mM, what is the plasma concentration of H2CO3? The pKa of carbonic acid is 3.57 . Plasma concentration of H2CO3 mM
Answer #1

According to hesselbach Hender sen and equetion Ka + loy [ Base] (acid] If pM <7 then acidic It pH>7 then basic If pH =7 thenp= pke + log [Ho,] (Hocog] 12m M 7.40= 357+ log (Haos] log CH, 6.3) 3.83 = 12 mn 3-8 3 (H,lo,] 12MM = 6760-82975387 (H, (03)